### v1

# Two-Step QAOA: Enhancing Quantum Optimization by Decomposing One-Hot Constraints in QUBO Formulations

Abstract
The Quantum Approximate Optimization Algorithm (QAOA) has shown promise in solving combinatorial optimization problems by leveraging quantum computational power. We propose a simple approach, the Two-Step QAOA, which aims to improve the effectiveness of QAOA by decomposing problems with one-hot encoding QUBO (Quadratic Unconstrained Binary Optimization) formulations. By identifying and separating the problem into two stages, we transform soft constraints into hard constraints, simplifying the generation of initial conditions and enabling more efficient optimization. The method is particularly beneficial for tackling complex societal problems that often involve intricate constraint structures.
1 Introduction
Quantum combinatorial optimization problems using quantum computers have seen various methods proposed for both problem formulation and actual solution implementation through the use of QAOA (Quantum Approximate Optimization Algorithm) [1]. QAOA can convert the formulation into operations called quantum gates according to theory and implement them on the quantum circuits of actual quantum computers or corresponding simulators. Currently, many societal problems worldwide are being tackled using algorithms like QAOA and quantum annealing, with applications expected in diverse fields such as logistics, finance, and natural language processing [2]. However, the performance of QAOA is highly dependent on the problem formulation. When dealing with QUBO (Quadratic Unconstrained Binary Optimization) formulations for societal problems [3], the relationship between constraint conditions and cost functions often leads to unnecessary parameter adjustments and computational load. Efficient computation necessitates innovations in formulation and subsequent embedding into quantum circuits.
When introducing penalty terms in QUBO, the adjustment variables that determine the influence of constraint conditions are introduced, making adjustments difficult depending on the state of the formulation. To eliminate the uncertainties associated with such soft constraints, the use of the Quantum Alternating Operator Ansatz (QAOA) is effective. This approach treats some soft constraints as hard constraints and solves these challenges by leveraging quantum entanglement and the mixer Hamiltonian.
In this paper, we propose a simple approach called "Two-Step QAOA." This approach enhances the effectiveness of QAOA by decomposing the QUBO formulation for constraints known as one-hot constraints and utilizing the Quantum Alternating Operator Ansatz. QUBO problems are widely used to generalize various combinatorial optimization problems and have been extensively studied due to their flexibility and expressiveness. However, QUBO problems often mix soft and hard constraints, complicating the adjustment process and making optimization challenging.
Two-Step QAOA divides specific problems into two quantum circuit stages, transforming soft constraints into hard constraints, simplifying the creation of initial conditions, and enabling more efficient optimization. In the first stage, we distinguish between soft constraints and cost functions from the structure of the QUBO formulation, handling the soft constraints separately and preparing the process to transform them into hard constraints and corresponding quantum states. In the second stage, we use the hard constraints transformed from the soft constraints to create a new QUBO formulation, optimizing it more effectively using QAOA.
The objective of this study is to use the Two-Step QAOA approach to reduce the complexity of setting initial conditions and enhance the overall performance of QAOA. This is expected to open new avenues for solving complex societal problems such as logistics, resource allocation, and scheduling. Additionally, since the quantum state generation process is heuristic, various improvements, including approximations, can be anticipated.
2 Quantum Adiabatic Process and Hamiltonians
The quantum adiabatic process is a method used in quantum computing and quantum mechanics to solve optimization problems. The fundamental principle of the adiabatic process is that a quantum system remains in its ground state if a given Hamiltonian changes slowly enough. This is known as the Adiabatic Theorem.
  1. 1.
    Initial Hamiltonian $H_{i}$: Start with an initial Hamiltonian whose ground state is easy to prepare. The system is initialized in this ground state.
  2. 2.
    Final Hamiltonian $H_{f}$: Define a final Hamiltonian whose ground state encodes the solution to the optimization problem.
  3. 3.
    Interpolating Hamiltonian: The system evolves according to a time-dependent Hamiltonian that slowly changes from the initial Hamiltonian to the final Hamiltonian. The interpolating Hamiltonian is given by:
    $H(t)=(1-s(t))H_{i}+s(t)H_{f}$ (1)
    where $s(t)$ is a smooth function that varies from 0 to 1 as time $t$ progresses from 0 to $T$.
  4. 4.
    Adiabatic Evolution: The system is allowed to evolve slowly enough so that it remains in its ground state throughout the process. By the end of the evolution (at time $T$), the system’s state will be close to the ground state of the final Hamiltonian $H_{f}$, providing the solution to the optimization problem.
In quantum computing, Hamiltonians are operators corresponding to the total energy of the system. They play a crucial role in defining the dynamics and evolution of quantum systems. For an optimization problem, the Hamiltonian is constructed such that its ground state (the state with the lowest energy) represents the optimal solution. QAOA can be utilized to solve this.
3 QAOA (Quantum Approximate Optimization Algorithm)
QAOA involves transforming a combinatorial optimization problem into a quantum bit (qubit) arrangement problem. The goal is to minimize a cost function by controlling the state of qubits, starting from a initial state where all qubits are equally likely to be in state 0 or 1.
QAOA consists of the following steps:
3.1 Problem Formulation
The combinatorial optimization problem is typically expressed as a cost function (objective function). This cost function is transformed into a Hamiltonian for qubits. The Hamiltonian $H_{C}$ corresponds to the problem’s cost function, and the goal is to minimize it.
3.2 Initialization
In a basic explanation of QAOA, we first assume a simple state, preparing all qubits in a uniform superposition. This state means that all qubits are equally likely to be in the state 0 or 1. The initial state is represented as follows:
$|\psi_{0}\rangle=\frac{1}{\sqrt{2^{n}}}\sum_{z\in\{0,1\}^{n}}|z\rangle$ (2)
3.3 Application of Cost and Mixer Hamiltonians
QAOA alternately applies the cost Hamiltonian $H_{C}$ and the mixer Hamiltonian $H_{M}$. These Hamiltonians determine the problem’s characteristics and the qubits’ exploration behavior.
  • Cost Hamiltonian $H_{C}$: This represents the problem’s cost function and reflects the energy landscape directly. It is applied as follows:
    $U(H_{C},\gamma)=e^{-i\gamma H_{C}}$ (3)
  • Mixer Hamiltonian $H_{M}$: This maintains the qubits in an exploratory state, preserving the superposition and preventing the qubits from getting trapped in local minima. It is typically applied as follows:
    $U(H_{M},\beta)=e^{-i\beta H_{M}}$ (4)
In each step of QAOA, these unitary operators are applied alternately, evolving the quantum state.
3.4 Parameter Optimization
The application of the cost and mixer Hamiltonians requires time parameters $\gamma$ and $\beta$. These parameters are optimized using classical optimization methods to maximize the probability of reaching the optimal solution.
3.5 Measurement and Result Extraction
The final state is measured to obtain the solution to the optimization problem. The measurement result is a bit string that represents the problem’s optimal solution.
3.6 Mathematical Representation
The overall process of QAOA is mathematically represented as follows. Starting from the initial state, the cost and mixer Hamiltonians are applied $p$ times:
$|\psi(\gamma,\beta)\rangle=U(H_{M},\beta_{p})U(H_{C},\gamma_{p})\cdots U(H_{M},\beta_{1})U(H_{C},\gamma_{1})|\psi_{0}\rangle$ (5)
The final state $|\psi(\gamma,\beta)\rangle$ is measured to find the optimal bit string. The parameters $\gamma$ and $\beta$ are chosen to minimize the following objective function:
$\langle\psi(\gamma,\beta)|H_{C}|\psi(\gamma,\beta)\rangle$ (6)
In general QAOA, the initial state is chosen to be a superposition of all qubits, and the mixer Hamiltonian is typically selected as $X$ Hamiltonian.
4 Quantum Alternating Operator Ansatz
The Quantum Alternating Operator Ansatz [4] is a generalized version of the Quantum Approximate Optimization Algorithm designed to solve combinatorial optimization problems. This approach allows for greater flexibility in selecting the operators used in the algorithm, which can be tailored to specific problems for improved performance. While the computational steps generally adhere to those of the basic QAOA, there are some changes in the initialization of the state and the setting of the mixer Hamiltonian.
4.1 Mixer Hamiltonian
In standard QAOA, the mixer Hamiltonian is typically the $X$ operator. However, in the Quantum Alternating Operator Ansatz, the mixer Hamiltonian is constructed using Pauli operators. This allows for explorations that cannot be achieved with just the $X$ Hamiltonian.
4.2 Initialization
Instead of assuming a simple state, the initial state in the Quantum Alternating Operator Ansatz is prepared as an eigenstate with complex quantum entanglement corresponding to the mixer Hamiltonian.
5 Cost and Constraint Terms in QAOA formulation
In QAOA, constructing the cost Hamiltonian involves creating terms that minimize the objective function and satisfy necessary constraints. These terms are balanced using tuning parameters, presenting several challenges during the process.
5.1 Formulating Cost and Constraint Terms
The cost Hamiltonian $H_{C}$ is composed of a cost term $C$, which represents the objective function, and a penalty term $P$, which represents the constraints that must be satisfied. These terms are connected using a tuning parameter $\lambda$:
$H_{C}=C+\lambda P$ (7)
5.2 Balancing the Terms
The primary challenge is to balance the cost and constraint terms to ensure that the solution minimizes the objective function while satisfying all constraints. If the tuning parameter $\lambda$ is too large, the solution might prioritize satisfying the constraints over minimizing the cost, leading to suboptimal solutions. Conversely, if $\lambda$ is too small, the constraints might be violated.
5.3 Iterative Tuning of Parameters
Determining the appropriate value for $\lambda$ often requires iterative optimization. This process involves adjusting $\lambda$ iteratively until a satisfactory balance is achieved, making sure that the final solution both minimizes the cost and satisfies the constraints.
The search space expands with the addition of constraint terms, making it harder to find the global minimum.
6 Incorporating Hard Constraints into Mixer Hamiltonian in QAOA: A Two-Step Initialization Approach
In specific formulations, such as when the constraint conditions can be described using an XY operator in the mixer Hamiltonian, it is possible to relax the constraints from the cost function. These constraints are excluded from the cost Hamiltonian as hard constraints and instead incorporated into the mixer Hamiltonian. This approach is particularly effective for problems involving one-hot encoding, where the constraints can naturally be expressed using the XY mixer. By leveraging the symmetry and properties of the XY operator, one-hot encoded variables can be efficiently managed, ensuring that only one variable in a set is activated while the others remain off. Utilizing the mixer Hamiltonian necessitates the preparation of an initial state corresponding to it. Various methods have been proposed for this purpose, including the use of Dicke states. In this case, the initial state can be easily created using a two-step approach.
6.1 Identify Constraints
Determine which constraints can be expressed using the XY operator in the mixer Hamiltonian. This involves analyzing the problem to identify constraints that can be represented in this form.
$H_{constraints}=P$ (8)
6.2 Construct Mixer Hamiltonian
In this problem, QAOA is executed twice, each time using a different mixer Hamiltonian.
In the first QAOA, the mixer Hamiltonian $H_{M_{1}}$ is
$H_{M_{1}}=H_{X}=X$ (9)
utilizing the X mixer.
Next, to incorporate the identified constraints, the mixer Hamiltonian is formulated as $H_{M_{2}}=H_{XY}$ using the XY operator. The mixer Hamiltonian will take the following form:
$H_{M_{2}}=H_{XY}=\sum_{i,j}XY_{ij}+\ldots$ (10)
where $XY_{ij}$ represents the XY interaction between qubits $i$ and $j$.
6.3 Modify Cost Hamiltonian
Exclude the identified constraints from the cost Hamiltonian $H_{C}$, treating them as hard constraints now embedded in the mixer Hamiltonian. The modified cost Hamiltonian will focus solely on the objective function.
$H_{C}=C$ (11)
6.4 Prepare Initial State
Prepare the initial state that corresponds to the mixer Hamiltonian. Various methods have been proposed for this purpose, including the use of Dicke states [5]. In this case, the initial state can be efficiently created using a two-step approach:
Start with an initial superposition state:
$|\psi_{0}\rangle=H^{\otimes n}|0\rangle^{\otimes n}$ (12)
Figure 1: Left: Conventional QAOA. Right: Improved Two-Step QAOA.
6.5 First QAOA Execution
Run QAOA with the constraints as the cost function and $X$ as the mixer Hamiltonian to generate the intermediate quantum state:
$|\psi_{\text{intermediate}}(\gamma_{1},\beta_{1})\rangle=U(H_{X},\beta_{1})U(H_{\text{constraints}},\gamma_{1})\cdots U(H_{X},\beta_{1})U(H_{\text{constraints}},\gamma_{1})|\psi_{0}\rangle$ (13)
6.6 Second QAOA Execution
Use the intermediate quantum state as the starting point and run QAOA with the objective function as the cost Hamiltonian and the XY mixer Hamiltonian:
$|\psi_{\text{final}}(\gamma_{2},\beta_{2})\rangle=U(H_{\text{XY}},\beta_{2})U(H_{C},\gamma_{2})\cdots U(H_{\text{XY}},\beta_{2})U(H_{C},\gamma_{2})|\psi_{\text{intermediate}}(\gamma_{1},\beta_{1})\rangle$ (14)
Optimize the parameters $\gamma_{1},\gamma_{2},\beta_{1},\beta_{2}$ through classical optimization techniques to find the optimal solution.
7 Discussion
The two-step approach in QAOA simplifies the handling of extraneous tuning parameters by focusing on distinct phases for constraint satisfaction and objective function minimization. This bifurcation into two steps, while adding complexity to the overall procedure, offers several advantages that could potentially enhance the efficiency and applicability of QAOA in solving complex Hamiltonian problems.
7.1 Advantages
By isolating the constraint satisfaction phase from the objective function minimization phase, the two-step approach reduces the number of tuning parameters that need simultaneous optimization. This separation allows for more straightforward tuning of each phase individually, thereby simplifying the parameter space and potentially improving convergence rates. Additionally, the initial state preparation using the X mixer Hamiltonian helps to embed the constraints directly into the quantum state, which can lead to more efficient state evolution during the subsequent QAOA steps.
7.2 Optimization and Efficiency
The complexity introduced by the two-step process can be managed by optimizing each step independently. By first running QAOA to generate an intermediate state that satisfies the constraints, we ensure that the subsequent optimization phase focuses solely on minimizing the objective function without being encumbered by constraint satisfaction. This targeted optimization in each phase can lead to better performance and more accurate results, as each step is tailored to specific aspects of the problem.
7.3 Future Applications
The flexibility of the two-step approach allows for the potential incorporation of advanced optimization techniques and hybrid classical-quantum algorithms. By leveraging the strengths of both quantum and classical optimization methods, this approach could be refined to handle larger and more complex problems in the future. The ability to adapt and optimize each phase of the QAOA process separately enhances the method’s robustness and applicability across a broader range of quantum computing problems.
Acknowledgments
I would like to express my gratitude to Takao Tomono from Keio University and independent researcher Shunsuke Sotobayashi for their invaluable contributions and support.